OpenStudy (anonymous):

Help please! Need help with Lagrange multipliers. Calc 3. f(x,y)=4x^3+y^2 ; constraint 2x^2 +y^2=1

OpenStudy (zarkon):

where are you stuck?

OpenStudy (anonymous):

I've found F_x, F_y, and F_lambda, but in order to solve for x or y you have to divide, which is where I am stuck. x or y can be zeros, so I;m not sure what to do with that. Also, the Ys seem to cancel when solving F_y. Hopefully that makes sense.

OpenStudy (zarkon):

what did you get for \(f_x,f_y\)

OpenStudy (anonymous):

Let Lambda = L. F_x=12x^-4Lx F_y=2y-2Ly

OpenStudy (anonymous):

12x^2**

OpenStudy (zarkon):

What are you using \(F(x,y,\lambda)\)?

OpenStudy (anonymous):

whoops

OpenStudy (anonymous):

Yes I am using F(x,y,L)

OpenStudy (anonymous):

I found the partials for each from F(x,y,L); then I set them equal to zero.

OpenStudy (zarkon):

ok..so you have one equation being \[2y=\lambda 2y\] correct?

OpenStudy (anonymous):

yes

OpenStudy (zarkon):

what does that tell you?

OpenStudy (anonymous):

L=1

OpenStudy (zarkon):

is that all?

OpenStudy (anonymous):

well, y=2Ly/2 can't really see much else

OpenStudy (zarkon):

does y=0 satisfy the equation too?

OpenStudy (zarkon):

I need to do something...be back in a few mins

OpenStudy (anonymous):

Okay

OpenStudy (zarkon):

back...you you see what I am talking about?

OpenStudy (anonymous):

No, I'm not quite sure what you mean.

OpenStudy (zarkon):

if \(2y=\lambda y\) then either \(y=0\) or \(y\ne0\) and therefroe \(\lambda=1\)

OpenStudy (zarkon):

forgot a 2 \(2y=\lambda 2y\)

OpenStudy (zarkon):

\[2\cdot 0=\lambda 2\cdot 0\] \[0=0\]

OpenStudy (anonymous):

Right, but we don't know what y actually equals, but if it did equal zero, does that not mean that we can not divide by it?

OpenStudy (zarkon):

if y=0 we cannot divide by it

OpenStudy (anonymous):

Right, that's where I am stuck. In order to solve for anything in the equation 2y=2Ly we have to divide.

OpenStudy (zarkon):

i'm just looking at cases...either y=0 or it doesn't

OpenStudy (zarkon):

if y=0 then it satisfies the equation \(2y=\lambda 2y\) if \(y\neq 0\) then we can divide by it and get \(\lambda=1\)

OpenStudy (anonymous):

What happens if y is 0?

OpenStudy (zarkon):

if y=0 what does x have to be?

OpenStudy (anonymous):

+- 1/sqrt2 ?

OpenStudy (zarkon):

correct...so you now have 2 critical points...are there others?

OpenStudy (zarkon):

what happens then if y is not zero and lambda=1

OpenStudy (anonymous):

3x=1?

OpenStudy (zarkon):

yep

OpenStudy (anonymous):

so what do I do for that?

OpenStudy (zarkon):

so what is x...and therefore what is y?

OpenStudy (anonymous):

x=1/3, y=sqrt7/3?

OpenStudy (zarkon):

so now you have two more critical numbers

OpenStudy (zarkon):

you forgot the \(\pm\)

OpenStudy (anonymous):

Hm, I just plug in to original and see which is biggest at this point, right?

OpenStudy (zarkon):

yes and smallest (i guess it depends on what the original question asked for )

OpenStudy (anonymous):

Max and min. Okay, I think I get it better now. Check the constraint if x or y is 0...that's where I was lost. Thanks for your help!!

OpenStudy (zarkon):

no problem

OpenStudy (zarkon):

you do have more critical points though

OpenStudy (zarkon):

one of your equations was \[12x^2=\lambda 4x\] \(x=0\) satisfies this equation

OpenStudy (anonymous):

So here x=L/3

OpenStudy (zarkon):

if \(x=0\) then \(y=\pm 1\)

OpenStudy (anonymous):

Oh okay. So that would be the last one right

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

So CPs: (0,+-1), (+- 1/sqrt2,0), (1/3, sqrt7/3) Right?

OpenStudy (anonymous):

scratch the 0 on the second one

OpenStudy (zarkon):

yes

OpenStudy (zarkon):

the last one is +-

OpenStudy (anonymous):

Oh yeah. So quick question: if my constraint were to be something like xy=1, then x and y both could not equal 0 right?

OpenStudy (zarkon):

correct

OpenStudy (anonymous):

Okay, then I fully understand it now! Appreciate the man.

OpenStudy (anonymous):

help*

OpenStudy (zarkon):

no problem